m/10-naoh ki ph value  nikal dijiye this is 11 class question

Grade:11

2 Answers

Arun
5 years ago
pH is negative log of hydrogen ion in a solution pH is represented by pH=-log(H+) it ranges from 1-14
here NaOH represent sodium hydoxide
and n/10 represent normality of the solution it can be writen as 0.1N
one mole of NaOH produces one sodium ion and one hydroxide ion 
so the pH can be calulated as follows:
ph=14-0.1(M)=13.0
 
 
Rajdeep
5 years ago
HEY THERE! 
 
The given solution is of Sodium Hydroxide, which is a Base, and hence furnishes Hydroxyl (OH-) ions.
So, we can easily calculate the pOH first from it’s given concentration (1/10 M) and then, find it’s pH from the formula:
pH + pOH = 14
 
Concentration = M/10
This means = 0.1 M = 10-1M
pOH = – log [OH-]
= – log (10-1)
= 1 (log 10)
= 1
 
Now, we get the pOH to be 1.
 
pH = 14 – pOH
= 14 – 1
= 13 (Ans)
 
Thanks!